OpenStudy (anonymous):

Find the points of inflection: y=x^3-8x^3+x

OpenStudy (anonymous):

find the second derivative of the function and set it equal to zero, this will give you the x value for the points of inflection

OpenStudy (anonymous):

So far i got the derivative, 4x^3 -24x^2 + 1

OpenStudy (anonymous):

ah yes sorry about that.. ^^"

OpenStudy (campbell_st):

find the 1st and 2nd derivatives solve the 1st derivative to find the stationary points find the 2nd derivative... test the stationary points f''(a) > 0 min f"(a) = 0 horizontal pt of inflexion f"(a) < 0 a maximum... if there isn't a horizontal point of inflexion then solve the 2nd derivative. then test there is a change in concavity either side of the solution. if there is then you have a point of inflexion.

OpenStudy (anonymous):

the equation is x^4-8x^3 + x

OpenStudy (anonymous):

@campbell_st oh so to find the Point of Inflection or (P.I.) I would have to find the 2nd Derivative?

OpenStudy (anonymous):

I thought to find P.I. i would only have to do First Derivative.. :o

OpenStudy (campbell_st):

yes... but you need to find the solutions to the 1st derivative as your point of inflexion may be a horizontal point... so the way I do it is find 1st derive then 2nd derive solve 1st derive = 0 test the solutions in the 2nd derive by substituting. then solve the 2nd derive... test for a change in concavity by picking a value close to the solution on either side... and seeing a change in the sign.

OpenStudy (campbell_st):

there are 2 types of inflexion points.... 1. horizontal... found at the 1st derive ann tested in the 2nd derive. 2. just point of inflexion... found as a solution to the 2nd derive and then test.

OpenStudy (campbell_st):

hope it make sense

OpenStudy (anonymous):

ah yes iremember now! :D

OpenStudy (anonymous):

yeah i was just too lazy to look over my notes lol xD

OpenStudy (campbell_st):

and just graphing your curve there is a horizontal pt of inflexion and a normal one... .

OpenStudy (anonymous):

@campbell_st can you help me with another one? ^^"

OpenStudy (campbell_st):

ok... f'(x) = 4x^3 - 24x^2 + 1 oops may have made a mistake... no obvious solutions there... f"(x) = 12x^2 - 48x solutions to f'(x) are x = 0 or 4 check change in concavity x = -1 f''(-1) = 60 x = 1 f''(1) = -36 so pt of inflexion at x = 0 tes x= 4 f''(3) = 108 - 144 = -36 f''(5) = 300 - 240 = 60 change of sign so pt of inflexion at x = 4

OpenStudy (anonymous):

yeah i got it now thx :)